Solve the following IVP

y"-10y'+9y=5t;y(0)=-1,y'(0)=2


1
Expert's answer
2021-07-14T14:41:07-0400

Given the differential equation,


"y''-10y'+9y=5t"

such that


"y(0)=-1,y'(0)=2"

The homogenous part of the equation is:


"y''-10y'+9y=0"

Let "m" be the root of the auxiliary equation such that:


"m^2-10m+9=0\\
(m-9)(m-1)=0\\
m =9 \text{ and } m=1"

are the roots of the equation.

Therefore the particular solution is:


"y_c = c_1e^t + c_2e^{9t}"

We determine the particular solution to the non-homogenous part of the given differential equation using the method of undetermined coefficient.


"y''-10y'+9y=5t"


The solution is of the form:


"y_p = a_1+a_2t"

Solving for the unknown constants "a_1" and "a_2" :


"y_p' = a_2\\
y_p'' = 0\\"

Substituting the particular solution into the differential equation, we get:


"-10a_2+9(a_1+a_2t) =5t\\"

Simplifying further, we have:


"9a_1-10a_2+9a_2t = 5t"

By equating coefficients:

"9a_1-10a_2=0\\
9a_2t = 5t"

Solving for the system yields:


"a_1 = \frac{50}{81}\\
a_2 = \frac{5}{9}"

Substituting "a_1 \text{ and } a_2 \text{ into } y_p, \text{ we have }:"


"y_p = \frac{5t}{9}+\frac{50}{81}"

Thus, the general solution is:


"y= y_c+y_p = c_1e^t + c_2e^{9t}+ \frac{5t}{9}+\frac{50}{81}"

We now solve for the unknown constants using the initial conditions:


"\dfrac{dy}{dt} = c_1e^t+9c_2e^{9t}+\frac{5}{9}"

Substituting y(0)=-1 into y (the general solution), we have:


"c_1+c_2+\frac{50}{81}=-1"

Substitute y'(0) = 2 into "\frac{dy}{dt}" we have:

"c_1+9c_2+\frac{5}{9}=2"

Solving the system gives:


"c_1 = -2\\
c_2 = \frac{31}{38}"

Substitute the values of "c_1 \text{ and } c_2" into the general solution, we have:


"y= \frac{1}{81}(31e^{9t}-162e^t+45t+50)"

which is the required solution.